OpenStudy (anonymous):

What are the solutions of x2 + 5x − 24 = 0? Show all work for full credit.

OpenStudy (anonymous):

Factoring! An important mathematical skill. x^2 + 5x - 24 = 0 Set up your two parentheses: (x _ _) (x _ _) = 0 We need numbers that sum to +5 and multiply together to -24. This requires us to know the factors of 24: (1,24),(2,12),(3,8),(4,6). Now, you'll notice that (3,8) has a difference of 5 (8 - 3 = 5), so let's use those! (x _ 3) (x _ 8) = 0 Now, we need this to produce x-terms which sum to +5, so 8 must be positive, and the 3 negative. (x - 3) (x + 8) = 0 Now, we simply set each term equal to 0 separately: x-3 = 0 --> x = 3 x+8 = 0 --> x = -8